# When LLMs Meet Cunning Questions: A Fallacy Understanding Benchmark for Large Language Models

### v1

@@ Appendix B Translations of Prompts set in FLUB
Our designed prompts without Chain-of-Thought for Task 1 and Task 2 are shown in Figure 5. The English translations of all our designed prompts for each task of FLUB are shown in Figure 6 and Figure 7.
Appendix C Automatic Evaluation Prompt in Task 3
@@
Appendix D Scoring Guidelines in Task 3
@@
For the LLMs’ output response to each question, please refer to the given annotated answer to score the output response. This score will be used to measure the quality of the model response and the degree of matching with the answer.
@@
D.1 Definition and Examples of Score 1
@@
The LLM’s response doesn’t point out the irrationality or humor in the given question, which is completely unrelated to the correct answer.
D.1.1 The First Example of Score 1
Question:
Since instant noodles can be eaten dry, then dry crispy noodles can also be boiled? (既然方便面可以干吃,那干脆面也可以水煮?)
@@ D.1.1 The First Example of Score 1
Yes, dry crispy noodles can also be boiled. Boil in boiling water until softened and serve. (是的,干脆面也可以水煮。将其放入沸水中煮至软化即可食用。)
@@
Dry crispy noodles are crispy and delicious ready-to-eat snack food. You can eat them as soon as you open the package. What people need is its crispness. So its taste will be lost after being boiled. Therefore, crispy noodles cannot be boiled. (干脆面是一种香脆可口的即食型休闲食品,打开包装就可以吃,吃的就是它的香脆,水煮之后会失去它的口感。所以干脆面不可以水煮。)
D.1.2 The Second Example of Score 1
Question:
To say that Ruozhi Bar is all a bunch of retards is that considered a compliment or an insult? (说弱智吧都是一帮弱智,算夸奖还是辱骂?)
@@ D.1.2 The Second Example of Score 1
This sentence is insulting because it uses the derogatory term “retarded” to describe a group of people and expresses contempt and insult to them. (这句话是辱骂,因为它使用贬义词汇“弱智”来形容一群人,表达了对他们的轻视和侮辱。)
@@
“Ruozhi Bar” is the name of the post bar, and the IQ of the members inside has nothing to do with what the post bar name is. “Retarded” means intelligence below the normal level, and to say ‘Ruozhi Bar” is a bunch of retarded is an insult. (“弱智吧”是贴吧名,里面成员的智商与贴吧名是什么无关。“弱智”的意思是智力低于正常水平,说弱智吧都是一帮弱智是一种辱骂。)
D.2 Definition and Examples of Score 2
@@
Although LLM’s response points out the unreasonable/humorous points in the question, the types and key points pointed out are inconsistent with the opinions of explanations.
D.2.1 The First Example of Score 2
Question:
Dreams are all reversed. I dream of myself looking in the mirror, so I am the one in the mirror. (梦都是反的,我梦到自己照镜子,所以我才是镜子里的人。)
@@ D.2.1 The First Example of Score 2
The humor of this sentence lies in the literal interpretation of the saying “Dreams are all reversed”. Usually, people think that the scene in a dream is the opposite of reality, but in this sentence, the author dreams of himself looking in the mirror, if we understand that “Dreams are all reversed”, then the author is actually a reflection in the mirror, which is obviously an absurd interpretation of reality, and therefore has a sense of humor. (这个句子的幽默之处在于对“梦都是反的”这一说法的字面解释。通常,人们认为梦中的情景与现实相反,但在这个句子中,作者梦到自己照镜子,如果按照“梦都是反的”来理解,那么作者实际上是镜子里的倒影,这显然是对现实的荒诞解释,因此具有幽默感。)
@@
“Dreams are all reversed” means that the dream is opposite to reality. If “I dream of myself looking in the mirror” is contrary to reality, the reality should be “I do not look in the mirror in reality”, rather than “I dream of myself looking in the mirror, and I am the person in the mirror”. (“梦都是反的”是指梦境与现实是相反的,“我梦到自己照镜子”与现实相反,应该是“现实中没有照镜子”,而不是“我梦到自己照镜子,反过来我是镜中人”。)
D.2.2 The Second Example of Score 2
Question:
If all three of Nezha’s heads suffered from cerebral thrombosis for ten years, did he have cerebral thrombosis for ten years or thirty years? (哪吒三头都得了十年脑血栓那是算十年脑血栓还是三十年脑血栓?)
@@ D.2.2 The Second Example of Score 2
Nezha is a fictional character in Chinese mythology that does not exist. Cerebral thrombosis is a cerebrovascular disease, usually caused by blood clotting in the blood vessels of the brain resulting in brain hypoxia and ischemia. Therefore, the claim that Nezha’s three heads have cerebral thrombosis is not realistic, let alone calculate whether it is ten years or thirty years. (哪吒是中国神话中的虚构人物,没有真实存在。脑血栓是一种脑血管疾病,通常是由于脑部血管内的血液凝固形成的血栓导致脑部缺氧缺血。因此,哪吒三头得脑血栓的说法是不符合实际的,更不用计算是十年还是三十年。)
@@
Nezha is an individual, assuming Nezha’s three heads all have cerebral thrombosis, it should be something that happens at the same time, and will not produce accumulation, so it should not be 30 years. (哪吒是一个个体,假设哪吒的三个头都得了脑血栓,那应该是同时发生的事情,并不会产生累加,所以不应该是三十年。)
D.3 Definition and Examples of Score 3
@@
The LLM’s response correctly points out the type of irrationality/humor, but the key points of irrationality/humor pointed out by the correct answer are inconsistent. Alternatively, the response is consistent with the key points of irrationality/humor pointed out by the correct answer, but the judgment on the type of question is inconsistent.
D.3.1 The First Example of Score 3
Question:
The drug is partially poisonous with side effects, so we should not take the drug in parts. (是药三分毒,所以药不能分三次喝。)
@@ D.3.1 The First Example of Score 3
The humor of this sentence lies in the misunderstanding of the literal and actual meaning of the idiom “The drug is partially poisonous with side effects”. This idiom is meant to say that any drug has potential side effects, while the sentence misunderstands “partially” to mean that the drug should not be taken in parts, and thus draws the absurd conclusion that the drug cannot be taken in three separate doses, which is obviously a comical distortion of the original meaning. (这句子的幽默之处在于对成语“是药三分毒”的字面意思和实际含义的误解。成语本意是说任何药物都有潜在的副作用,而句子却把“三分”误解为药物应该分成三个部分服用,因此得出了一个荒谬的结论,即药物不能按照三次分开服用,这显然是对原意的一种滑稽曲解。)
@@
“The drug is partially poisonous with side effects” refers to any drug will have certain side effects if used improperly. “Partially” indicates that to a certain extent, there is no meaning of “in parts”. Therefore, the conclusion “we should not take the drug in parts” cannot be drawn from this sentence. (“是药三分毒”指的是任何的药品如果使用不当的话都会存在一定的副作用,“三分”表示的是在一定程度上,并没有分三次的意思,所以“是药三分毒”这句话并不能得出”药不能分三次喝“这一结论。)
D.3.2 The Second Example of Score 3
Question:
Did a knee-jerk reflex test on a friend, and the result was that his leg never lifted again after the hammer was knocked down. This reflex arc is too long! (给朋友做膝跳测试,结果锤下去之后他的腿再也没抬起过,这反射弧也太长了吧!)
@@ D.3.2 The Second Example of Score 3
The knee-jerk reflex test is a simple test to check the function of the nervous system. Normally, a light tap on the tendon below the knee will cause the leg to make a quick leg lift motion. This sentence humorously exaggerates the amount of time the friend does not respond, indicating that the reflex arc is unusually long, when in fact, if there is really no response, it may indicate a neurological problem rather than the reflex arc length. In addition, using “never lifted again” has a more humorous effect because it indicates a permanent state, which is impossible in reality. (膝跳反射测试是一种检查神经系统功能的简单测试,正常情况,轻轻敲击膝盖下方的腱会导致腿部做出快速的抬腿动作。这句话幽默地夸张了朋友没有反应的时间,暗示反射弧异常地长,实际上如果真的没有反应,那可能表明存在神经系统问题,而不是反射弧长。此外,用“再也没抬起过”增加了幽默效果,因为它暗示了一种永久性的状态,这在现实中是不可能的。)
@@
“His leg never lifted again after the hammer” indicates that the friend’s leg was broken, and is not related to the length of the reflex arc, here is actually a humorous way to express the matter of “breaking the friend’s leg”. (“锤下去之后他的腿再也没抬起过”说明把朋友的腿打断了,与反射弧长短没有关系,这里实际上是用幽默的方式将“把朋友腿打断”表达出来。)
D.4 Definition and Examples of Score 4
@@
The response correctly points out the type of irrationality/humor in the question, and the key points pointed out are correct, but the interpretation is not consistent with the correct answer.
D.4.1 The First Example of Score 4
Question:
In a vegetable market, one finds vegetative persons instead. So it must have been a job market before. (去素材库发现几个植物人躺在里面,原来之前这里是人材市场。)
@@ D.4.1 The First Example of Score 4
The humor of this sentence lies in the pun on the word “vegetative”. Usually, “vegetative” refers to patients who have fallen into a persistent vegetative state due to a major illness or injury, but here it is used to refer to “vegetable” in the literal sense, i.e., real plants. At the same time, “job market” usually refers to a market for hiring “people” with the appropriate skill, but here it is used teasingly to refer to a market with vegetative “people”. This linguistic coincidence creates an absurd humorous effect. (这个句子的幽默之处在于使用了“植物人”这个词语的双关意味。通常,“植物人”指的是那些因为重大疾病或伤害而陷入持续植物状态的病人,但在这里它被用来指代字面意义上的“植物”,即真正的植物。同时,“人材市场”通常是指招聘人才的市场,而在这里它被戏谑地用来指一个植物市场,这种语言上的巧合创造了一种荒谬的幽默效果。)
@@
The vegetable market refers to the market where vegetables are sold. “Vegetative” is a special human state similar to the living state of plants, and the human body cannot lie in the vegetable market. (素材库是指销售蔬菜的市场。“植物人”是与植物生存状态相似的特殊的人体状态,人体是无法躺进素材库中的。)
D.4.2 The Second Example of Score 4
Question:
If you drink 82 Lafite, then you’ll live at least 82 years. (喝了82年拉菲至少能活82岁。)
@@ D.4.2 The Second Example of Score 4
The humor or irrationality of the sentence is that it establishes a false cause-and-effect relationship, implying that drinking a certain vintage of wine guarantees living to the same age as the vintage, which obviously has no scientific basis. In fact, a person’s lifespan is affected by many factors, including genetics, lifestyle, medical conditions, and so on, rather than a single diet. In addition, this statement also carries a kind of humorous exaggeration, because it uses an absurd logic to attract attention and increase the interest of the words. (该句子的幽默或不合理之处在于它建立了一种错误的因果关系,暗示喝了某年份的葡萄酒能保证活到与年份相同的年龄,这显然没有科学依据。实际上,人的寿命受多种因素影响,包括遗传、生活方式、医疗条件等,而非单一的饮食习惯。此外,这种说法也带有一种幽默的夸张,因为它用一种荒谬的逻辑来吸引注意,增加话语的趣味性。)
@@
“82 Lafite” is a network buzzword that originated from scenes of opening Lafite wine in movies and TV series. Since the Lafite red wine produced in 1982 is of good quality and expensive, “82 Lafite” is synonymous with high standards to set off the character’s extravagant spending and taste, rather than “Lafite has been drunk for 82 years”. (82年的拉菲,网络流行语,源自影视剧中开拉菲的桥段,由于82年的拉菲红酒品质好、价格昂贵,所以“82年的拉菲”是高规格的代名词,来衬托人物的挥金如土和品味,而不是“喝了八十二年的拉菲”的意思。)
D.5 Definition and Examples of Score 5

### v2

@@ Appendix A Our Designed Prompts for FLUB
Our designed prompts without Chain-of-Thought for FLUB are shown in Figure 4. And the Chain-of-Thought prompts for Task 1 and Task 2 are presented in Figure 5. For the convenience of reading, we provide both prompts and their English translations. Note that we use Chinese prompts in specific experiments because our data samples are in Chinese.
Appendix B Automatic Evaluation Prompt in Task 3
@@
Appendix C Scoring Guidelines in Task 3
@@
For the LLMs’ output response to each input cunning text, please refer to the given annotated answer to score the output response. This score will be used to measure the quality of the model response and the degree of matching with the answer.
@@
C.1 Definition and Examples of Score 1
@@
The LLM’s response doesn’t point out the irrationality or humor in the given text, which is completely unrelated to the correct answer.
C.1.1 The First Example of Score 1
Input Text:
Since instant noodles can be eaten dry, then dry crispy noodles can also be boiled? ({CJK*}UTF8gbsn既然方便面可以干吃,那干脆面也可以水煮?)
@@ C.1.1 The First Example of Score 1
Yes, dry crispy noodles can also be boiled. Boil in boiling water until softened and serve. ({CJK*}UTF8gbsn是的,干脆面也可以水煮。将其放入沸水中煮至软化即可食用。)
@@
Dry crispy noodles are crispy and delicious ready-to-eat snack food. You can eat them as soon as you open the package. What people need is its crispness. So its taste will be lost after being boiled. Therefore, crispy noodles cannot be boiled. ({CJK*}UTF8gbsn干脆面是一种香脆可口的即食型休闲食品,打开包装就可以吃,吃的就是它的香脆,水煮之后会失去它的口感。所以干脆面不可以水煮。)
C.1.2 The Second Example of Score 1
Input Text:
To say that Ruozhi Bar is all a bunch of retards is that considered a compliment or an insult? ({CJK*}UTF8gbsn说弱智吧都是一帮弱智,算夸奖还是辱骂?)
@@ C.1.2 The Second Example of Score 1
This sentence is insulting because it uses the derogatory term “retarded” to describe a group of people and expresses contempt and insult to them. ({CJK*}UTF8gbsn这句话是辱骂,因为它使用贬义词汇“弱智”来形容一群人,表达了对他们的轻视和侮辱。)
@@
“Ruozhi Bar” is the name of the post bar, and the IQ of the members inside has nothing to do with what the post bar name is. “Retarded” means intelligence below the normal level, and to say ‘Ruozhi Bar” is a bunch of retarded is an insult. ({CJK*}UTF8gbsn“弱智吧”是贴吧名,里面成员的智商与贴吧名是什么无关。“弱智”的意思是智力低于正常水平,说弱智吧都是一帮弱智是一种辱骂。)
C.2 Definition and Examples of Score 2
@@
Although LLM’s response points out the unreasonable/humorous points in the given text, the types and key points pointed out are inconsistent with the opinions of explanations.
C.2.1 The First Example of Score 2
Input Text:
Dreams are all reversed. I dream of myself looking in the mirror, so I am the one in the mirror. ({CJK*}UTF8gbsn梦都是反的,我梦到自己照镜子,所以我才是镜子里的人。)
@@ C.2.1 The First Example of Score 2
The humor of this sentence lies in the literal interpretation of the saying “Dreams are all reversed”. Usually, people think that the scene in a dream is the opposite of reality, but in this sentence, the author dreams of himself looking in the mirror, if we understand that “Dreams are all reversed”, then the author is actually a reflection in the mirror, which is obviously an absurd interpretation of reality, and therefore has a sense of humor. ({CJK*}UTF8gbsn这个句子的幽默之处在于对“梦都是反的”这一说法的字面解释。通常,人们认为梦中的情景与现实相反,但在这个句子中,作者梦到自己照镜子,如果按照“梦都是反的”来理解,那么作者实际上是镜子里的倒影,这显然是对现实的荒诞解释,因此具有幽默感。)
@@
“Dreams are all reversed” means that the dream is opposite to reality. If “I dream of myself looking in the mirror” is contrary to reality, the reality should be “I do not look in the mirror in reality”, rather than “I dream of myself looking in the mirror, and I am the person in the mirror”. ({CJK*}UTF8gbsn“梦都是反的”是指梦境与现实是相反的,“我梦到自己照镜子”与现实相反,应该是“现实中没有照镜子”,而不是“我梦到自己照镜子,反过来我是镜中人”。)
C.2.2 The Second Example of Score 2
Input Text:
If all three of Nezha’s heads suffered from cerebral thrombosis for ten years, did he have cerebral thrombosis for ten years or thirty years? ({CJK*}UTF8gbsn哪吒三头都得了十年脑血栓那是算十年脑血栓还是三十年脑血栓?)
@@ C.2.2 The Second Example of Score 2
Nezha is a fictional character in Chinese mythology that does not exist. Cerebral thrombosis is a cerebrovascular disease, usually caused by blood clotting in the blood vessels of the brain resulting in brain hypoxia and ischemia. Therefore, the claim that Nezha’s three heads have cerebral thrombosis is not realistic, let alone calculate whether it is ten years or thirty years. ({CJK*}UTF8gbsn哪吒是中国神话中的虚构人物,没有真实存在。脑血栓是一种脑血管疾病,通常是由于脑部血管内的血液凝固形成的血栓导致脑部缺氧缺血。因此,哪吒三头得脑血栓的说法是不符合实际的,更不用计算是十年还是三十年。)
@@
Nezha is an individual, assuming Nezha’s three heads all have cerebral thrombosis, it should be something that happens at the same time, and will not produce accumulation, so it should not be 30 years. ({CJK*}UTF8gbsn哪吒是一个个体,假设哪吒的三个头都得了脑血栓,那应该是同时发生的事情,并不会产生累加,所以不应该是三十年。)
C.3 Definition and Examples of Score 3
@@
The LLM’s response correctly points out the type of irrationality/humor, but the key points of irrationality/humor pointed out by the correct answer are inconsistent. Alternatively, the response is consistent with the key points of irrationality/humor pointed out by the correct answer, but the judgment on the type of cunning text is inconsistent.
C.3.1 The First Example of Score 3
Input Text:
The drug is partially poisonous with side effects, so we should not take the drug in parts. ({CJK*}UTF8gbsn是药三分毒,所以药不能分三次喝。)
@@ C.3.1 The First Example of Score 3
The humor of this sentence lies in the misunderstanding of the literal and actual meaning of the idiom “The drug is partially poisonous with side effects”. This idiom is meant to say that any drug has potential side effects, while the sentence misunderstands “partially” to mean that the drug should not be taken in parts, and thus draws the absurd conclusion that the drug cannot be taken in three separate doses, which is obviously a comical distortion of the original meaning. ({CJK*}UTF8gbsn这句子的幽默之处在于对成语“是药三分毒”的字面意思和实际含义的误解。成语本意是说任何药物都有潜在的副作用,而句子却把“三分”误解为药物应该分成三个部分服用,因此得出了一个荒谬的结论,即药物不能按照三次分开服用,这显然是对原意的一种滑稽曲解。)
@@
“The drug is partially poisonous with side effects” refers to any drug will have certain side effects if used improperly. “Partially” indicates that to a certain extent, there is no meaning of “in parts”. Therefore, the conclusion “we should not take the drug in parts” cannot be drawn from this sentence. ({CJK*}UTF8gbsn“是药三分毒”指的是任何的药品如果使用不当的话都会存在一定的副作用,“三分”表示的是在一定程度上,并没有分三次的意思,所以“是药三分毒”这句话并不能得出”药不能分三次喝“这一结论。)
C.3.2 The Second Example of Score 3
Input Text:
Did a knee-jerk reflex test on a friend, and the result was that his leg never lifted again after the hammer was knocked down. This reflex arc is too long! ({CJK*}UTF8gbsn给朋友做膝跳测试,结果锤下去之后他的腿再也没抬起过,这反射弧也太长了吧!)
@@ C.3.2 The Second Example of Score 3
The knee-jerk reflex test is a simple test to check the function of the nervous system. Normally, a light tap on the tendon below the knee will cause the leg to make a quick leg lift motion. This sentence humorously exaggerates the amount of time the friend does not respond, indicating that the reflex arc is unusually long, when in fact, if there is really no response, it may indicate a neurological problem rather than the reflex arc length. In addition, using “never lifted again” has a more humorous effect because it indicates a permanent state, which is impossible in reality. ({CJK*}UTF8gbsn膝跳反射测试是一种检查神经系统功能的简单测试,正常情况,轻轻敲击膝盖下方的腱会导致腿部做出快速的抬腿动作。这句话幽默地夸张了朋友没有反应的时间,暗示反射弧异常地长,实际上如果真的没有反应,那可能表明存在神经系统问题,而不是反射弧长。此外,用“再也没抬起过”增加了幽默效果,因为它暗示了一种永久性的状态,这在现实中是不可能的。)
@@
“His leg never lifted again after the hammer” indicates that the friend’s leg was broken, and is not related to the length of the reflex arc, here is actually a humorous way to express the matter of “breaking the friend’s leg”. ({CJK*}UTF8gbsn“锤下去之后他的腿再也没抬起过”说明把朋友的腿打断了,与反射弧长短没有关系,这里实际上是用幽默的方式将“把朋友腿打断”表达出来。)
C.4 Definition and Examples of Score 4
@@
The response correctly points out the type of irrationality/humor in the input text, and the key points pointed out are correct, but the interpretation is not consistent with the correct answer.
C.4.1 The First Example of Score 4
Input Text:
In a vegetable market, one finds vegetative persons instead. So it must have been a job market before. ({CJK*}UTF8gbsn去素材库发现几个植物人躺在里面,原来之前这里是人材市场。)
@@ C.4.1 The First Example of Score 4
The humor of this sentence lies in the pun on the word “vegetative”. Usually, “vegetative” refers to patients who have fallen into a persistent vegetative state due to a major illness or injury, but here it is used to refer to “vegetable” in the literal sense, i.e., real plants. At the same time, “job market” usually refers to a market for hiring “people” with the appropriate skill, but here it is used teasingly to refer to a market with vegetative “people”. This linguistic coincidence creates an absurd humorous effect. ({CJK*}UTF8gbsn这个句子的幽默之处在于使用了“植物人”这个词语的双关意味。通常,“植物人”指的是那些因为重大疾病或伤害而陷入持续植物状态的病人,但在这里它被用来指代字面意义上的“植物”,即真正的植物。同时,“人材市场”通常是指招聘人才的市场,而在这里它被戏谑地用来指一个植物市场,这种语言上的巧合创造了一种荒谬的幽默效果。)
@@
The vegetable market refers to the market where vegetables are sold. “Vegetative” is a special human state similar to the living state of plants, and the human body cannot lie in the vegetable market. ({CJK*}UTF8gbsn素材库是指销售蔬菜的市场。“植物人”是与植物生存状态相似的特殊的人体状态,人体是无法躺进素材库中的。)
C.4.2 The Second Example of Score 4
Input Text:
If you drink 82 Lafite, then you’ll live at least 82 years. ({CJK*}UTF8gbsn喝了82年拉菲至少能活82岁。)
@@ C.4.2 The Second Example of Score 4
The humor or irrationality of the sentence is that it establishes a false cause-and-effect relationship, implying that drinking a certain vintage of wine guarantees living to the same age as the vintage, which obviously has no scientific basis. In fact, a person’s lifespan is affected by many factors, including genetics, lifestyle, medical conditions, and so on, rather than a single diet. In addition, this statement also carries a kind of humorous exaggeration, because it uses an absurd logic to attract attention and increase the interest of the words. ({CJK*}UTF8gbsn该句子的幽默或不合理之处在于它建立了一种错误的因果关系,暗示喝了某年份的葡萄酒能保证活到与年份相同的年龄,这显然没有科学依据。实际上,人的寿命受多种因素影响,包括遗传、生活方式、医疗条件等,而非单一的饮食习惯。此外,这种说法也带有一种幽默的夸张,因为它用一种荒谬的逻辑来吸引注意,增加话语的趣味性。)
@@
“82 Lafite” is a network buzzword that originated from scenes of opening Lafite wine in movies and TV series. Since the Lafite red wine produced in 1982 is of good quality and expensive, “82 Lafite” is synonymous with high standards to set off the character’s extravagant spending and taste, rather than “Lafite has been drunk for 82 years”. ({CJK*}UTF8gbsn82年的拉菲,网络流行语,源自影视剧中开拉菲的桥段,由于82年的拉菲红酒品质好、价格昂贵,所以“82年的拉菲”是高规格的代名词,来衬托人物的挥金如土和品味,而不是“喝了八十二年的拉菲”的意思。)
C.5 Definition and Examples of Score 5